OpenStudy (anonymous):

sovle (y-3)-(y+8)=9y

OpenStudy (saifoo.khan):

(y-3)-(y+8)=9y y - 3 - y - 8 = 9y Shuffle, y - y - 9y = + 8 + 3 -9y = 11 y = -11/9

OpenStudy (anonymous):

thanks appreciate the help